Shift report 20011106 16:00-24:00

From: Apache (apache@pii3.brahms.bnl.gov)
Date: Tue Nov 06 2001 - 23:54:28 EST

  • Next message: Djamel Ouerdane: "H2 Pid"

    ------------------------------------------------------------
    Date: Tue Nov  6 23:54:27 2001, Shift: 16:00-24:00
    Supervisor: rw, with: 
    Log book#10 pages 106 - 108
    
    Shift Summary:
    ------------------------------------------------------------
    1. Beam from the previous shift was dumped at 18:30. Finally 
    at 23:45 beam was cogged. Blue/yellow=24.5/17.8 ZDC=700.
    
    2. FV and HI worked for checkimg the noise of the drift chambers,
    by changing the anode HV. 
    	run 	trigger 6     Comments
    	5693                  drift chamber diagonostic
     	5694		      same	
    	5695		      same
    	5696		      same
    	5697        2K        FS 20deg-1/10A  MRS 90deg-4A
    			      a very short run, just before 
    			      dumping.
            5701    just started
    
    3. No new problems.
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b30 : Tue Nov 06 2001 - 23:55:09 EST